OpenStudy (anonymous):

how would i solve: Find a real number, t, such that the line containing the points (8, 9) and (t, 4) has a slope of 3

OpenStudy (anonymous):

use point slope formula y-9=3(x-8) [the 9 and 8 are from point (8,9) and the 3 is from the slope simplify to get y=3x+15 Then plug in 4 for y and solve for the x

OpenStudy (anonymous):

I did and I got -11/3=x, but when I type it into my homework it says it's wrong?

OpenStudy (precal):

You should use y=3x-15 by the way you should use the slope formula (9-4)/(8-t)=3 cross multiply 3(8-t)=5 finish it and you will get the correct t value.